# NSmark: Null Space Based Black-box Watermarking Defense Framework for Pre-trained Language Models

### v1

@@ Abstract
Pre-trained language models (PLMs) have emerged as critical intellectual property (IP) assets that necessitate protection. Although various watermarking strategies have been proposed, they remain vulnerable to Linear Functionality Equivalence Attack (LFEA), which can invalidate most existing white-box watermarks without prior knowledge of the watermarking scheme or training data. This paper further analyzes and extends the attack scenarios of LFEA to the commonly employed black-box settings for PLMs by considering Last-Layer outputs (dubbed LL-LFEA). We discover that the null space of the output matrix remains invariant against LL-LFEA attacks. Based on this finding, we propose NSmark, a task-agnostic, black-box watermarking scheme capable of resisting LL-LFEA attacks. NSmark consists of three phases: (i) watermark generation using the digital signature of the owner, enhanced by spread spectrum modulation for increased robustness; (ii) watermark embedding through an output mapping extractor that preserves PLM performance while maximizing watermark capacity; (iii) watermark verification, assessed by extraction rate and null space conformity. Extensive experiments on both pre-training and downstream tasks confirm the effectiveness, reliability, fidelity, and robustness of our approach. Code is available at https://github.com/dongdongzhaoUP/NSmark.
NSmark: Null Space Based Black-box Watermarking Defense Framework for Pre-trained Language Models
Haodong Zhao1,2, Jinming Hu1, Peixuan Li1, Fangqi Li1, Jinrui Sha1, Peixuan Chen2, Zhuosheng Zhang1,∗, Gongshen Liu1,thanks: Corresponding authors. 1Shanghai Jiao Tong University 2Tencent 1 {zhaohaodong, hujinming, peixuan.li, solour_lfq, sjrgckym, zhangzs, lgshen}@sjtu.edu.cn 2 noahchen@tencent.com
@@ 1 Introduction
During past decades, with the rapid development of artificial intelligence (AI), pre-trained language models (PLMs) have achieved superior performance and been applied in a wide range of fields. At the same time, training high-performance PLMs requires a large amount of data and computing resources, thus PLMs can be regarded as valuable intellectual property (IP). With the deployment and application of machine learning as a service (MLaaS) platforms, companies sell well-trained PLMs as commodities. Once sold models are illegally stolen, distributed or resold, the rights of the model owner will be seriously infringed. Therefore, it is necessary to protect IP of PLMs.
Figure 1: Illustration of different watermark schemes against LFEA/LL-LFEA.
@@
Watermarking techniques have been widely applied for protecting IP of deep learning models Chen et al. (2024); He et al. (2024). By incorporating identifiable information, these methods serve as a means to verify model ownership and provide proof of authenticity. Existing schemes can be divided into white-box and black-box categories according to whether the model parameters need to be accessed in verification. Among them, black-box schemes are more applicable when model parameters are inaccessible in most real-life scenarios, such as deployment as API.
@@
However, protecting the IP of PLMs through watermarking presents significant challenges. Since PLMs can be deployed for various downstream tasks post-training Zhang et al. (2023), it is crucial for watermark schemes to be task-independent. Besides, recent studies have revealed vulnerabilities and shortcomings in existing watermarking techniques Li et al. (2023a). The proposed Linear Functionality Equivalence Attack (LFEA) is easy to conduct and can invalidate most existing white-box watermarks without knowledge of the watermarking scheme or the training data by exploiting linear invariance. Considering Last-Layer outputs, we further analyze and expand the attack scenarios black-box settings using model outputs (dubbed LL-LFEA).
@@
To address these problems, we first explore the characteristics of the model output. We find that the null space of the matrix composed of the output vectors of model is invariant under LL-LFEA. Based on this, we propose a new null space verification method that can resist LL-LFEA attack. This method uses a new metric, the Null Space Matching Degree (NSMD). NSMD measures the degree of match between the output matrix of the suspicious model and the null space of the protected PLM. Finally, we propose NSmark, a null space based task-agnostic black-box watermarking scheme for PLMs. NSmark  uses identity information to generate all watermark-related elements, and uses WER and NSMD to verify the watermark. Spread spectrum modulation technology and an extra extractor are also introduced to enhance the watermark performance.
@@
(i) We analyze the threat of LFEA on output-based watermark, and propose LL-LFEA, which can destroy the watermark embedded in the output vector without affecting the performance of downstream tasks.
(ii) We find that the null space of the matrix composed of the output vectors of model is invariant under LL-LFEA and thus propose a new null space verification method NSmark  that can resist LL-LFEA. Notably, NSmark  is task-agnostic that uses both new null space verification and signature verification to resist LL-LFEA.
(iii) We conduct comprehensive experiments by applying NSmark  to various models of both pre-training and downstream tasks. Experimental results demonstrate the effectiveness, fidelity, reliability and robustness of NSmark.
@@ 2 Related Work
Watermarking for PLMs. With the rise of pre-training in NLP, recent work has explored watermarking specific to PLMs. BadPre Jia et al. (2022) introduces a task-agnostic backdoor attack only for MLM-based PLMs. Hufu Xu et al. (2024) introduces a modality-agnostic approach for pre-trained Transformer models using the permutation equivariance property. Explanation as a Watermark Shao et al. (2024) addresses the limitations of backdoor-based techniques by embedding multi-bit watermarks into feature attributions using explainable AI. Shen et al. (2021); Zhang et al. (2023) propose task-agnostic backdoor attacks by assigning high-dimensional vectors as trigger set labels, but their effectiveness is sensitive to downstream classifier initialization. Wang and Kerschbaum (2021) introduced an auxiliary neural network for watermark embedding using weights from the main network. Wu et al. (2022) proposed a task-agnostic embedding loss function, but didn’t consider the need for triggers to reflect the model owner’s identity. Cong et al. (2022) introduced a black-box watermarking scheme for PLMs, but its applicability is limited due to the discrete nature of word tokens. Unfortunately, these schemes are vulnerable to attacks by LFEA or LL-LFEA in principle.
Watermark Removal Attacks. DNN watermarking faces various removal attempts. Common methods include fine-tuning Adi et al. (2018) and pruning Han et al. (2015). Fine-pruning Liu et al. (2018) combines these approaches for increased effectiveness. Knowledge Distillation Hinton (2015) techniques can also inadvertently remove watermarks while reducing model size. Lukas et al. (2022) propose a new attack method called Neuron Reordering to swap neurons within the same hidden layer of a DNN to disrupt embedded watermarks in the model’s parameters. Li et al. (2023a) introduce a powerful LFEA for white-box watermarks, applying linear transformations to model parameters, effectively destroying embedded watermarks while preserving the model’s original functionality. Fraud attacks include overwriting Wang and Kerschbaum (2019) and ambiguity attacks Zhu et al. (2020) also pose a great threat to watermarks.
@@ 3.1 Threat Model
In white-box watermarking schemes, high-dimensional model parameters are often used as watermark information. For PLMs, since the output of the last layer is a high-dimensional vector, we can use a method similar to the white-box watermarking scheme to embed watermarks in the output. However, embedding identity information into the high-dimensional output vector will face the threat of LFEA-like attacks, which is proposed to destroy watermark information embedded in model parameters by linearly transforming parameters of intermediate layers. Next, we discuss the specific form of linear isomorphism attacks in this scenario.
Assuming that the attacker knows the watermark information is embedded in the output of the PLM, and hopes to destroy the watermark at a very low attack cost (without modifying the model structure or fine-tuning the model) without affecting the normal task performance of the model. As shown in Figure 2, we give an attack method that can meet this requirement and provide a proof below. Output vector $\vec{x}$ is obtained from PLM, which is also the input of downstream model. After a series of linear and nonlinear layers, the prediction result $y$ is obtained. The attacker attempts to destroy the watermark by changing the output vector of PLM, but does not affect the final prediction result of the model, that is, the attacker hopes to change $\vec{x}$ to $\varphi(x)$ and input it into the downstream network, and corresponding prediction result $y^{\prime}=y$. The sufficient condition for this result is that the modification of the PLM output vector can be offset after passing through the first linear layer of the downstream network. Let the parameter of the first linear layer of the downstream network be $W$, then the attacker hopes $W^{\prime}\varphi(\vec{x})=W\vec{x}$, then we have $\varphi(\vec{x})=W^{\prime,\dagger}W\vec{x}=Q\vec{x}$, where $Q=W^{\prime,\dagger}W$ and $W^{\prime,\dagger}$ is the pseudo-inverse of $W^{\prime}$ Li et al. (2023a). In order not to lose information in linear transformation (otherwise downstream tasks will be affected uncontrollably), $Q$ must be a reversible matrix. This proves that attacker can perform a linear transformation on $\vec{x}$ to destroy the watermark embedded in the output vector without affecting the performance of downstream tasks. We call this attack the Last-Layer Linear Functionality Equivalence Attack (LL-LFEA).
Figure 2: The schematic diagram of model inference flow before and after LL-LFEA attack.
@@ 3.2 Null Space Verification Theory
LL-LFEA performs linear transformation on output vector of PLM and can destroy the watermark embedded in it, which means previous watermark verification methods will be significantly affected. We find that the null space of the matrix composed of the output vector is invariant under the LL-LFEA attack, and propose to use the null space matching degree to verify whether the model is embedded with watermarks.
Theorem 1.
Before and after LL-LFEA, the null space of the output matrix of PLM remains unchanged for the same input set. (Proofs in Appendix A.1)
Figure 3: The overall workflow of NSmark.
@@ 3.3 Null Space Match Degree (NSMD)
We define NSMD by introducing the distribution of elements in a matrix, which is obtained by matrix multiplication of the output matrix $A$ of any PLM without watermark and the null space matrix $N$ of $f_{wm}$. In $H_{(n\times p)}=A_{(n\times m)}\times N_{(m\times p)}$, $H_{i,j}=\alpha_{i}\cdot\beta_{j}$ is the dot product of the i-th row vector of $A$ and the j-th column vector of $N$. We define NSMD of $A$ and $N$ as:
@@ 3.4 Overall Framework of NSmark
As shown in Figure 3, NSmark  scheme includes three modules: watermark generation, watermark embedding, and watermark verification:
@@ 3.4.1 Watermark Generation
The workflow of watermark generation is shown in Algorithm 1. We hope that the generated watermark contains the owner’s identity information. First, the digital signature $sig=\textbf{Sign}(m)$ is generated from the identity information message $m$. In order to ensure that the trigger $t$ has a unique mapping relationship with the $sig$, only one trigger is used. We use the trigger generation algorithm $\textbf{Encode}(\cdot)$ introduced in Li et al. (2023b) to obtain $t=\textbf{Encode}(sig,n=1)$. $t$ is inserted into clean sample $x$ of dataset $D$ to form a trigger set $D_{T}$.
In order to defend against ambiguous attacks, the verification trigger set $D_{V}$ used for null space verification also needs to be generated based on $sig$. A candidate pool $D_{NS}$ for generating null space verification data sets needs to be published, and then a fixed number of samples are selected from the $D_{NS}$ based on the digital signature as the verification data set $D_{V}$. We define the verification data set selection algorithm as $\textbf{Select}(sig)\rightarrow D_{V}$, which must be a deterministic algorithm, that is, for the same input, there must be the same output. In addition, we hope that the algorithm has different outputs for different inputs. Therefore, we choose a hash function and use a one-way hash chain to generate $D_{V}$. We hope that the index repetition rate obtained by different hash value mappings is low, so we hope that the data set $D_{NS}$ is as large as possible. The specific process of the $\textbf{Select}(\cdot)$ algorithm is shown in Algorithm 2.
In order to improve the robustness of the watermark, we introduce spread spectrum modulation technology as Feng and Zhang (2020). Spread spectrum modulation technology uses redundant bits to represent the original information. Figure 4 shows an example of 3×3$\times$3 × spreading. Please refer to the Appendix A.5.1 for the specific process $\textbf{SM}(sig)\rightarrow sig_{wm}$.
Figure 4: Example diagram of spread spectrum modulation.
@@ 3.4.2 Watermark Embedding
Before the training starts, make a copy of $f_{wm}$ as the frozen reference model $f_{ref}$. Then use the clean data set $D$ and the trigger set $D_{T}$ to train $f_{wm}$ and the extractor $E$. When taking $\{D,D_{T}\}$ as input, $f_{wm}$ will output $\{V,V^{T}\}$ and $f_{ref}$ will output $\{V_{ref},V_{ref}^{T}\}$, respectively. For $V^{T}$, $E$ maps it to obtain the signature $sig_{wm}$, and for $\{V,V_{ref},V_{ref}^{T}\}$, $E$ maps them to random vectors. After the training is completed, $f_{wm}$ is embedded with watermark. Then using $D_{V}$ as input, the output vectors are concatenated into a matrix $A$, and the corresponding null space matrix $N$ of $A$ is calculated as part of the key.
Three networks are involved in watermark embedding: the model $f_{wm}$ to be embedded with watermark, the reference model $f_{ref}$, and the extractor model $E$. Compared with directly embedding $sig$ into the output vector of $f_{wm}$, adding $E$ to map can reduce the side effect of watermark on the original performance. Watermark capacity is increased at the same time. We use mean square error loss (MSE) and similarity function $\operatorname{sim}$ to implement the above training process:
$L_{\text{match}}=\frac{1}{|D_{T}|}\sum_{x\in D_{T}}\operatorname{MSE}\left(E\left(f_{wm}(x)\right),{sig}_{sm}\right),$ (2)
$L_{\text{random }}$ $=\frac{1}{|D|}\sum_{x\in D}\operatorname{sim}\left(E\left(f_{wm}(x)\right),sig_{sm}\right)^{2}$ (3)
$+\frac{1}{|D_{T}|}\sum_{x\in D_{T}}\operatorname{sim}\left(E\left(f_{{ref}}(x)\right),sig_{sm}\right)^{2}$
$+\frac{1}{|D|}\sum_{x\in D}\operatorname{sim}\left(E\left(f_{{ref}}(x)\right),sig_{sm}\right)^{2}.$
We use cosine similarity as the $\operatorname{sim}$ function. The complete loss function of $E$ is $L_{\text{Extractor}}=\lambda_{1}L_{\text{match}}+(1-\lambda_{1})L_{\text{random}}$. When training $E$ using the loss, only the parameters of $E$ are trainable. The loss of $f_{wm}$ also consists of two parts: $L_{wm}=\lambda_{2}L_{match}+(1-\lambda_{2})L_{0}$. The content of this $L_{match}$ is the same as $L_{match}$ of $E$, but only the parameters of $f_{wm}$ are updated at this time, and $L_{0}$ is the original training loss function of PLM. During training, $E$ and $f_{wm}$ are trained alternately.

### v2

# NSmark: Null Space Based Black-box Watermarking Defense Framework for Language Models

@@ Abstract
Language models (LMs) have emerged as critical intellectual property (IP) assets that necessitate protection. Although various watermarking strategies have been proposed, they remain vulnerable to Linear Functionality Equivalence Attack (LFEA), which can invalidate most existing white-box watermarks without prior knowledge of the watermarking scheme or training data. This paper analyzes and extends the attack scenarios of LFEA to the commonly employed black-box settings for LMs by considering Last-Layer outputs (dubbed LL-LFEA). We discover that the null space of the output matrix remains invariant against LL-LFEA attacks. Based on this finding, we propose NSmark, a black-box watermarking scheme that is task-agnostic and capable of resisting LL-LFEA attacks. NSmark consists of three phases: (i) watermark generation using the digital signature of the owner, enhanced by spread spectrum modulation for increased robustness; (ii) watermark embedding through an output mapping extractor that preserves the LM performance while maximizing watermark capacity; (iii) watermark verification, assessed by extraction rate and null space conformity. Extensive experiments on both pre-training and downstream tasks confirm the effectiveness, scalability, reliability, fidelity, and robustness of our approach. Code is available at https://github.com/dongdongzhaoUP/NSmark.
Machine Learning, ICML
@@ 1 Introduction
Over the past few decades, language models (LMs) have achieved exceptional performance and found applications across a wide range of fields (Yu et al., 2025). However, training high-performance LMs requires vast amounts of data and significant computational resources, making these models valuable intellectual property (IP). With the rise of machine learning as a service (MLaaS) platforms, companies sell well-trained LMs as commodities and release APIs for public access. Once these models are illegally stolen, distributed or resold, the rights of the model owners are severely violated. Therefore, protecting the intellectual property of LMs is essential.
Figure 1: Illustration of different watermark schemes against LFEA/LL-LFEA. LFEA disables parameters based white-box schemes (Li et al., 2023a) and LL-LFEA disables output based black-box schemes (Section 3.1). NSmark  is secure against LL-LFEA using null space invariance.
@@
Watermarking techniques have been widely used to protect the IP of deep learning models (Chen et al., 2024; He et al., 2024; Carlini et al., 2024; Feng et al., 2024). By incorporating identifiable information, these techniques could verify model ownership and provide proof of authenticity. Existing watermarking schemes can be categorized into white-box and black-box approaches, depending on whether the model parameters need to be accessed in verification. Among these, black-box schemes are more applicable in real-life scenarios, where model parameters are often inaccessible, such as in cases where models are deployed as APIs.
@@
However, protecting the IP of LMs through watermarking presents significant challenges. Since LMs can be deployed for various post-training downstream tasks (Zhang et al., 2023), it is crucial that watermark schemes remain task independent. Furthermore, recent studies have revealed vulnerabilities and shortcomings in existing watermarking techniques (Li et al., 2023a). Specifically, the proposed Linear Functionality Equivalence Attack (LFEA) is simple to conduct and can compromise most existing white-box watermarks by exploiting linear invariance without knowledge of the watermarking scheme or the training data. As the hidden states and outputs of the last layer in LM are widely used for classification and generation tasks, we consider them, analyze and expand LFEA scenarios to black-box settings utilizing model outputs (dubbed LL-LFEA).
@@
In this work, we first explore the characteristics of the model output. We discover that the null space of the matrix composed of the model output vectors is invariant under LL-LFEA. Based on this finding, we propose a new null space verification method that can withstand the LL-LFEA attack. This method uses a new metric, the Null Space Matching Degree (NSMD). NSMD measures the degree of match between the output matrix of the suspicious model and the null space of the protected LM. Finally, we propose NSmark, a null-space-based task-agnostic black-box watermarking scheme for LMs. NSmark  uses identity information to generate all elements related to the watermark and uses the Watermark Extracting Rate (WER) and NSMD to verify the watermark, thus can pass through as shown in Figure 1. Spread spectrum modulation technology and an extra extractor are also introduced to enhance watermark performance.
@@
(i) We analyze the threat of LFEA on output-based watermark and propose LL-LFEA, which can destroy the watermark embedded in the output vector without affecting the performance of downstream tasks.
(ii) We find that the null space of the matrix composed of the output vectors of the model is invariant under LL-LFEA and thus propose a new null space verification method NSmark  which can resist LL-LFEA. Notably, NSmark  is task-agnostic that uses both new null space verification and signature verification to resist LL-LFEA.
(iii) We conduct comprehensive experiments by applying NSmark  to various models of pre-training and downstream tasks. The experimental results demonstrate the effectiveness, fidelity, reliability, and robustness of NSmark.
@@ 2 Related Work
Watermarking for LMs. With the rise of pre-training in NLP, recent work has explored watermarking specific to LMs. BadPre (Jia et al., 2022) introduced a task-agnostic backdoor attack only for MLM-based LMs. Hufu (Xu et al., 2024) introduced a modality-agnostic approach for pre-trained Transformer models using the permutation equivariance property. Explanation as a Watermark (Shao et al., 2024) addressed the limitations of backdoor-based techniques by embedding multi-bit watermarks into feature attributions using explainable AI. (Peng et al., 2023; Shetty et al., 2024) proposed Embeddings-as-a-Service (EaaS) watermarks to protect the intellectual property of EaaS providers. (Shen et al., 2021; Zhang et al., 2023) proposed task-agnostic backdoor attacks by assigning high-dimensional vectors as trigger set labels, but their effectiveness is sensitive to downstream classifier initialization. (Wang & Kerschbaum, 2021) introduced an auxiliary neural network for watermark embedding using weights from the main network. (Wu et al., 2022) proposed a task-agnostic embedding loss function, but didn’t consider the need for triggers to reflect the model owner’s identity. (Cong et al., 2022) introduced a black-box watermarking scheme for PLMs, but its applicability is limited due to the discrete nature of word tokens. Unfortunately, these schemes are vulnerable to attacks by LFEA or LL-LFEA in principle.
Watermark Removal Attacks. DNN watermarking faces various removal attempts. Common methods include fine-tuning (Adi et al., 2018) and pruning (Han et al., 2015). Fine-pruning (Liu et al., 2018) combines these approaches for greater effectiveness. Knowledge Distillation (Hinton, 2015) techniques can also inadvertently remove watermarks while reducing model size. (Shetty et al., 2024) show that existing EaaS watermarks can be removed by paraphrasing attack. (Lukas et al., 2022) propose a new attack method called Neuron Reordering to swap neurons within the same hidden layer of a DNN to disrupt embedded watermarks in the model’s parameters. (Li et al., 2023a) introduce a powerful LFEA for white-box watermarks, applying linear transformations to model parameters, effectively destroying embedded watermarks while preserving the model’s original functionality. Fraud attacks include overwriting (Wang & Kerschbaum, 2019) and ambiguity attacks (Zhu et al., 2020) also pose a great threat to watermarks.
Figure 2: The schematic diagram of model inference flow before and after LL-LFEA attack. LL-LFEA transforms the LM output and performs an inverse transform in the subsequent linear layer, leaving the final prediction unchanged.
@@ 3.1 Threat Model
In white-box watermarking schemes, high-dimensional model parameters are often used as watermark information. For LMs, since the output of the last layer is also high-dimensional, we can use a method similar to the white-box schemes to embed watermarks in the output. However, embedding identity information into the high-dimensional output vector will face the threat of LFEA-like attacks, which is proposed to destroy watermark information embedded in model parameters by linearly transforming parameters of intermediate layers. Next, we discuss the specific form of linear isomorphism attacks in this scenario.
Assume that the attacker knows that the watermark information is embedded in the LM output and seeks to remove the watermark with minimal attack cost (without modifying the model structure or fine-tuning the model) while ensuring that the model’s normal task performance remains unaffected. As shown in Figure 2, we propose an attack method that satisfies this requirement and provide a proof below.
The output vector $\vec{x}$ is generated by the LM and serves as input to the downstream model. After passing through a series of linear and non-linear layers, the prediction result $y$ is obtained. The attacker attempts to modify the output vector of the LM to destroy the watermark while ensuring that the final prediction remains unaffected. Specifically, the attacker changes $\vec{x}$ to $\varphi(x)$ and inputs it into the downstream model, so that the resulting prediction $y^{\prime}$ remains equal to the original prediction $y$.
The sufficient condition for this result is that the modification to the LM output vector is compensated for after passing through the first linear layer of the downstream network. Let the parameter matrix of the first linear layer in the downstream network be denoted by $W$. In this case, the attacker aims to satisfy the following condition: $W^{\prime}\varphi(\vec{x})=W\vec{x},$ which leads to $\varphi(\vec{x})=W^{\prime,\dagger}W\vec{x}=Q\vec{x},$ where $Q=W^{\prime,\dagger}W$ and $W^{\prime,\dagger}$ is the pseudo-inverse of $W^{\prime}$ (Li et al., 2023a). To avoid loss of information during the linear transformation (since this would adversely affect downstream tasks), $Q$ must be a reversible matrix. We present a simple method and analysis on how to quickly generate high-dimensional Q in Appendix A.1.
We show that the attacker can apply a linear transformation to $\vec{x}$ thereby destroying the watermark embedded in the output vector, while leaving the downstream task performance unchanged. We refer to this attack as the Last-Layer Linear Functionality Equivalence Attack (LL-LFEA). In addition to the theoretical analysis, the effectiveness of LL-LFEA is experimentally verified in Appendix B.1.
@@ 3.2 Null Space Verification Theory
LL-LFEA applies a linear transformation to the output vector of LM and can destroy the embedded watermark. As a result, previous watermark verification methods may be significantly impacted. We observe that the null space of the matrix composed of the output vector is invariant under the LL-LFEA attack. Based on this, we propose to use the null space matching degree to verify whether the model is embedded with watermarks.
Theorem 3.1.
Before and after LL-LFEA, the null space of the output matrix of LM remains unchanged for the same input set.
Proof.
See details in Appendix A.2. ∎
Figure 3: The overall workflow of NSmark. (i) In watermark generation, identity information is used generate $sig$. (ii) In watermark embedding, watermarked model $f_{wm}$ and extractor $E$ are trained with the participation of the reference model $f_{ref}$. (iii) In watermark verification, WER and NSMD collaborate to verify the identity of the model.
@@ 3.3 Null Space Match Degree (NSMD)
We define NSMD by introducing the distribution of elements in a matrix, which is obtained by multiplication of the matrix of the output matrix $A$ of any LM without watermark and the null space matrix $N$ of $f_{wm}$. In $H_{(n\times p)}=A_{(n\times m)}\times N_{(m\times p)}$, $H_{i,j}=\alpha_{i}\cdot\beta_{j}$ is the dot product of the $i$th row vector of $A$ and the $j$th column vector of $N$. We define NSMD of $A$ and $N$ as:
@@ 3.4 Overall Framework of NSmark
NSmark  includes three modules: watermark generation, watermark embedding, and watermark verification, as shown in Figure 3. We describe the modules as follows.
@@ 3.4.1 Watermark Generation
Algorithm 1 shows the watermark generation workflow. We hope that the generated watermark contains the owner’s identity information. First, the digital signature $sig=\textbf{Sign}(m)$ is generated from the identity information message $m$. To ensure that the trigger $t$ has a unique mapping relationship with $sig$, only one trigger is used. We use the trigger generation algorithm $\textbf{Encode}(\cdot)$ introduced in (Li et al., 2023b) to obtain $t=\textbf{Encode}(sig,n=1)$. $t$ is inserted into clean sample $x$ of dataset $D$ to form a trigger set $D_{T}$.
To defend against ambiguous attacks, the verification trigger set $D_{V}$ used for the null space verification also needs to be generated based on $sig$. A candidate pool $D_{NS}$ to generate null space verification data sets should be published, and then a fixed number of samples are selected from the $D_{NS}$ based on the digital signature as the verification data set $D_{V}$. We define the verification data set selection algorithm as $\textbf{Select}(sig)\rightarrow D_{V}$, which must be a deterministic algorithm, that is, for the same input, there must be the same output. In addition, we hope that the algorithm will have different outputs for different inputs. Therefore, we choose a hash function and use a one-way hash chain to generate $D_{V}$. We hope that the index repetition rate obtained by different hash-value mappings is low, so we hope that the data set $D_{NS}$ is as large as possible. The specific process of the $\textbf{Select}(\cdot)$ algorithm is shown in Algorithm 2.
To improve the robustness of the watermark, we introduce the spread spectrum modulation technology as (Feng & Zhang, 2020). Spread spectrum modulation technology uses redundant bits to represent the original information. Figure 4 shows an example of the spreading of 3×3$\times$3 ×. Please refer to Appendix A.6.1 for the specific process $\textbf{SM}(sig)\rightarrow sig_{wm}$.
Figure 4: Example diagram of spread spectrum modulation. Repeat $sig$ to obtain $sig_{repeat}$, then use $sm$ to modulate $sig_{repeat}$ to obtain the spread spectrum modulated digital signature $sig_{sm}$.
@@ 3.4.2 Watermark Embedding
Before the training starts, make a copy of $f_{wm}$ as the frozen reference model $f_{ref}$. Then use the clean data set $D$ and the trigger set $D_{T}$ to train $f_{wm}$ and the extractor $E$. When taking $\{D,D_{T}\}$ as input, $f_{wm}$ will output $\{V,V^{T}\}$ and $f_{ref}$ will output $\{V_{ref},V_{ref}^{T}\}$, respectively. For $V^{T}$, $E$ maps it to obtain the signature $sig_{wm}$, and for $\{V,V_{ref},V_{ref}^{T}\}$, $E$ maps them to random vectors. After the training is completed, $f_{wm}$ is embedded with the watermark. Then using $D_{V}$ as input, the output vectors are concatenated into a matrix $A$, and the corresponding null space matrix $N$ of $A$ is calculated as part of the key.
Three networks are involved in watermark embedding: the model $f_{wm}$ to be embedded with the watermark, the reference model $f_{ref}$ and the extractor model $E$. Compared to directly embedding $sig$ in the output vector of $f_{wm}$, adding $E$ to the map can reduce the side effect of the watermark on the original performance. The watermark capacity is increased at the same time. We use the mean square error loss (MSE) and the similarity function $\operatorname{sim}$ to implement the above training process:
$L_{\text{match}}=\frac{1}{|D_{T}|}\sum_{x\in D_{T}}\operatorname{MSE}\left(E\left(V^{T}\right),{sig}_{sm}\right),$ (2)
$L_{\text{random }}$ $=\frac{1}{|D|}\sum_{x\in D}\operatorname{sim}\left(E\left(V\right),sig_{sm}\right)^{2}$ (3)
$+\frac{1}{|D_{T}|}\sum_{x\in D_{T}}\operatorname{sim}\left(E\left(V_{ref}^{T}\right),sig_{sm}\right)^{2}$
$+\frac{1}{|D|}\sum_{x\in D}\operatorname{sim}\left(E\left(V_{ref}\right),sig_{sm}\right)^{2}.$
We use cosine similarity as the $\operatorname{sim}$ function. The complete loss function of $E$ is $L_{\text{Extractor}}=\lambda_{1}L_{\text{match}}+(1-\lambda_{1})L_{\text{random}}$. During training, only the parameters of $E$ are trainable. The loss of $f_{wm}$ also consists of two parts: $L_{wm}=\lambda_{2}L_{match}+(1-\lambda_{2})L_{0}$. The content of this $L_{match}$ is the same as $L_{match}$ of $E$, but only the parameters of $f_{wm}$ are updated at this time, and $L_{0}$ is the original LM training loss function. During training, $E$ and $f_{wm}$ are trained alternately.